OpenStudy (anonymous):

write an equivalent expression using radical notiation (a^4b^4)^1/5

OpenStudy (anonymous):

\[ \sqrt[5]{b^{4} a^{4}}\]

jimthompson5910 (jim_thompson5910):

\[\large (a^4b^4)^{\frac{1}{5}}=\sqrt[5]{a^4b^4}\]

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

no prblm